2. (10 pts.) Suppose particle A is 65 em from particle B and that particle A has a charge of 6 μC and that particle B has a charge of 11 pC. If k is 9 x 10 Nm/C2, what is the force between the two particles?
Answer #1

Here ,

for the force between two particles


force between two particles = k * q1 * q2/d^2

force between two particles = 9 *10^9 * 6 *10^-6 * 11 *10^-12/.65^2

force between two particles = 1.41 *10^-6 N

the force between two particles is 1.41 *10^-6 N